Suppose that the length of a rectangle is 3 inches longer than the width and that the perimeter of the rectangle is 78.
a) Set up an equation involving only L, the length of the rectangle.
Answer:
b) Solve this equation algebraically to find the length of the rectangle. Find the width as well.
Answer: Length ______, Width ______
Show your work here:

Answer by checkley71(8403)

L=W+3 OR W=L-3
2L+2W=78
2L+2(L-3)=78
2L+2L-6=78
4L=78+6
4L=84
L=84/4
L=21 ANSWER FOR THE LENGTH.
W=21-3
W=18 ANSWER FOR THE WIDTH.
PROOF
2*21=2+18=78
42+36=78
78=78